OpenStudy (anonymous):

Solve –x – y = 8 using the elimination method. -2x – y = -1

OpenStudy (raden):

directly subtract both equation, what you get ....

OpenStudy (raden):

–x – y = 8 -2x – y = -1 ------------- (-) ?????????????